A math professor purchased a condominium, for $60,000 in 1977. The real estate agent said the value would increase at an annual rate of 7% each year. Assuming the real estate agent was correct, in what year would her condo be worth ten times what she paid for it?
The interest rat is given and so is the increase but, how do I set up an equation to find the year?

Answer by stanbon(75887)

A math professor purchased a condominium, for $60,000 in 1977. The real estate agent said the value would increase at an annual rate of 7% each year. Assuming the real estate agent was correct, in what year would her condo be worth ten times what she paid for it?
-----
A(t) = 60,000*1,07^t
---------
10*60,0000 = 60,000*1.07^t
----
1.07^t = 10
-------
t = log(10)/log(1.07)
t = 1/0.0294
t = 34.03 years
------
Ans: 1977+34 = 2011
-------------------------
Cheers,
Stan H.
========================